OpenStudy (anonymous):

Six kilograms of an inferior tea is mixed with 3 kilograms of tea that costs $2 a kilogram more. The total price of the mixture is $24. What is the price of the inferior tea?

OpenStudy (jhannybean):

"Six kilograms of an inferior tea is mixed with 3 kilograms of tea that costs $2 a kilogram more" Let x = inferior tea, y = more expensive tea 6x + 3y = 2 "The total price of the mixture is $24." x + y = 24

OpenStudy (jhannybean):

Now that you have a system of equations, you can use the substitution or elimination method to solve for both variables.

OpenStudy (anonymous):

Can you expand it a bit further?

OpenStudy (jhannybean):

expand? What are we expanding.

OpenStudy (anonymous):

Is it 24$ per kilogram for the new mixture or is it total for all 9 kilograms? if it is total you can just use 6x + 3(x+2) = 24 and solve for x which gives you x = 2. If it is 24$ per kilogram it will be 6x + 3(x+2) = 9*24 which gives you x = 23.33

OpenStudy (shinalcantara):

Six kilograms of an inferior tea is mixed with 3 kilograms of tea that costs $2 a kilogram more. The total price of the mixture is $24 ---------------------------- Let 'x' be the price of inferior tea 6x + 3(x+2) = 24 Distribute 3 to the quantity (x+2) 6x + 3x + 6 = 24 transpose 6 to the right side 6x + 3x = 24 - 6 Combine like terms 9x = 18 Divide the whole equation by 9 x = 2 ---------------------------- Therefore the inferior tea costs $2 per kg